

********************************************************************
UARS WEEKLY STATUS REPORT - 09 JUN 2000
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
ACRIM II data, beginning with the inception of science data operations of
the ACRIM III experiment on 4/05/00, are being analyzed for
comparisons and validation of ACRIM III. The signatures of two large
amplitude 'sunspot deficits' (~ -0.15% of TSI), centered near 4/26/00 and
5/19/00, were observed by both instruments. These coincided with two
successive rotations of active regions with large sunspot area.
The ACRIM II data processing software has been rewritten using the
improvements employed in the approach developed for the ACRIM III
experiment. Reprocessing has begun and should be completed by the end
of Summer, 2000.
***********
PEM
***********
The PEM spectrometer daylight operation was interrupted beginning June 5
due to the power negative condition which existed with the UARS batteries.
The FOT will attempt to restore PEM daylight operations beginning June 9
by PMON enabling. The PEM spectrometer operation in daylight is active
under PMON control* when PMON is enabled.
At the present time, PEM energy deposition is not being generated.
Recent data configuration changes have affected PEM software causing multiple
failures. These failures are being reviewed by the PEM team. In order to
correct for new telemetry requirements, PEM will need to update processing
code with a patch delivery. Updated HEPS calibration parameters were
finalized and are ready for delivery. Updated MEPS quality calibration files
were finalized and are ready for delivery. Level 3 software errors,
discovered by Dr. Dean Pesnell, have been fixed; his suggestions have been
incorporated with in the PEM production code. Both updated HEPS and MEPS
calibration data, along with the Level 3 code changes, will influence the
in-situ energy deposition predictions, so new versions of PEM energy
deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Following the yaw maneuver on 5 June, TMONs 1 and 2 tripped
and turned off ACRIM, SOLSTICE, SUSIM, HRDI and PEM. HRDI
and PEM were operating in their day-only mode. The MPG
assisted the off-line engineer and the FOT in analyzing
the problem and scheduling instrument restarts.
The MPG attended the UARS Monthly Review Meeting held
on 8 June at GSFC. Processing of instrument Daily Activity
Plans continues with no problems. We are continuing to
assist the Project Scientist in scheduling of instrument
operations.
***********
CDHF
***********
CDHF notified all PI Software Representatives that UARS DAY 3182 was
re-ingested as Version 3 for better quality and more data.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
1 day SEN (5.0) PROD (1R-1C=0L)
1 day ATT (7.3) PROD (re-ingested L0) (1R-1C=0L)
1 day OBC (3.4) PROD (re-ingested L0) (1R-1C=0L)
1 day SEN (5.0) PROD (re-ingested L0) (1R-1C=0L)
1 day SUS (7.2) PROD (re-ingested L0) (1R-1C=0L)
The CDHF will load April 2000 current day DAAC data in June 2000.
UARS CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, HRDI,
MLS, SUSIM, Solstice, and WINDII.
- Conducting system tests..NRT, LO ingest ,production processing,
correlative ingest, DAAC transfers etc...
- Preparing to test GPI, ISAMS V12.0, and PEM V3.32 software
on the Rehost System.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 May 2000 (Orbit 47598, GMT Day 148) through 02 June 2000
(Orbit 47702, GMT Day 154).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 148/0025 47597 AXIS Off
WINDII 151/1052 47649 Off
HALOE 154/0736 47692 Off
The beta angle ranged between the angles of 39.7 degrees to 11.7 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+0.0 degrees on 05 June 2000 (DOY 157).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.2 degrees C. Additional battery performance data for each
battery on 26 May 2000 and 02 June 2000 are provided in Appendix D.
The clock error ranged from 11.6 to -13.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 00
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 18
% HALOE = 01
% PEM = 01
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/153 show the observatory data loss to be 295 hours, 38
minutes, 10.561 seconds (an increase of 1 minutes 43.480 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
The increase of 1 minutes 43.480 seconds occurred over one (1) day:
- 1 minutes 43.480 seconds on Day 149 due to a late AOS.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to test the new TMONs on the simulator. Once the TMONs have
been thoroughly tested, they will be presented for review and approval before
uplink.
The FOT has begun writing the online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure will
be tested on the simulator and put online for immediate execution should an
anomalous condition take place.
MEETINGS
The FOT attended the UARS Science Team Meeting held at the Washington D.C.
Convention Center on 31 May 2000. The FOT presented the current status of the
major subsystems as well as current projects and plans for the future.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for June 05, 2000 (DOY 157). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 June 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47597 148/0025 00-235 PEM PEM AXIS OFF
47605 148/1232 00-233 NBTR NBTR NOOPS D148
47616 149/0612 00-236 NBTR NBTR NOOPS D149
47623 149/1811 00-237 NBTR NBTR NOOPS D150
47628 150/0209 00-238 OBC Clk Adj 0.00 to 20.6 msec/day
47631 150/0615 00-237 NBTR NBTR NOOPS D150
47638 150/1814 00-239 NBTR NBTR NOOPS D151
47644 151/0353 00-240 OBC Clk Adj 20.6 to 0.00 msec/day
47646 151/0618 00-239 NBTR NBTR NOOPS D151
47649 151/1052 00-241 WINDII Power OFF
47653 151/1816 00-242 NBTR NBTR NOOPS D152
47661 152/0622 00-242 NBTR NBTR NOOPS D152
47661 152/0622 00-243 OBC Clk Adj 0.00 to -20.6 msec/day
47664 152/1054 00-242 NBTR NBTR NOOPS D152
47676 153/0547 00-244 NBTR NBTR NOOPS D153
47678 153/0954 00-245 OBC Clk Adj -20.6 to 0.00 msec/day
47683 153/1748 00-244 NBTR NBTR NOOPS D-153
47683 153/1821 00-246 NBTR NBTR NOOPS D-154
47692 154/0736 00-247 HALOE Power OFF - S/C LOW PWR
47694 154/1140 00-248 OBC Clk Adj fr 0.00 to 20.6 msec/day
47695 154/1320 00-246 NBTR NBTR NOOPS
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 June 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 June 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47628 150/02:24 (28 May 00) from 0.00 to 20.6 msec/day
47644 151/04:12 (30 May 00) from 20.6 to 0.00 msec/day
47661 152/06:22 (31 May 00) from 0.00 to -20.6 msec/day
47678 153/09:39 (01 June 00) from -20.6 to 0.00 msec/day
47694 154/11:43 (02 June 00) from 0.00 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 June 2000
26 May 2000 - GMT Day 147 Beta = 39.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------- ------ ------
1 +16.8/+0.0 -3.05/-4.31 +0.0/+0.0 15.7 0.000 0.0
2 +39.2/-33.6 +5.79/+4.51 +28.4/-15.2 24.6 1.030 21.8
3 +22.4/-44.8 +5.15/+3.55 +27.6/-16.0 24.6 1.028 21.6
02 June 2000 - GMT Day 154 Beta = 11.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ----------- ------- ------- ------
1 +22.4/+0.0 -3.05/-4.62 +0.0/+0.0 15.7 0.000 0.0
2 +39.2/-22.4 +6.44/+3.87 +30.4/-14.4 24.5 1.040 20.4
3 +22.4/-67.2 +5.15/+2.92 +29.6/-14.4 24.3 1.038 19.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 June 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-021 FOT 47597 148/0025 14170 Generic Late Acq.
00-021 FOT 47620 149/1228 14170 Generic Late Acq.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-007 FOT see list May 00 14170 Generic Late Acq.
SEVNTFW00/21